OpenStudy (anonymous):

Jason had to submit five animation projects as per his syllabus in the final semesters. He scored 88, 79, 76, and 90 points out of 100 in four projects. What should be his score in the fifth project so that the average of all the projects is at least 83?

OpenStudy (anonymous):

If you take the fraction (88+79+76+90+x)/5 = 83 and solve for x you will get what his grade would have to be.

OpenStudy (anonymous):

so (333+x)/5 = 83 or 333+x=415 so x = 82